thank u
5. The values of a variable X are: 1,3,9,5, 4, 2 Compute the following: x3 {X? [(x - 2) XS
Answer #1

Values of X are : 13.9.5, 4, 2 Ex = 7+3+9+5+4+2 | Ex= 30 O Ex* = (30) (Ewa 900 Ex==(7+ (3) +(9)+(5) 7 (4)? +(2) 2 Ex² = 49+9+